a car weighing 100kg is on a slope yhat makes an angle 30 degrees with the horizomtal ..the component of cars weight parallel to t
he slope is  

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
The mass of car=100 kg
so weight of car=100*10=1000 N (taking g=10 ms^-2)
theta=30 degrees
component of car’s weight parallel to slope
=mg*sin theta
=1000*sin 30
=1000*1/2
=500 N [Ans]
 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)